# Magnetosphere Evolution and Precursor-Driven Electromagnetic Signals in Merging Binary Neutron Stars

### v1

@@ 3.3.3 Pair Production and Gamma-ray Attenuation
This magnetic pair production is particularly significant in environments where the magnetic field strength approaches or exceeds the critical value, $B_{c}=4.414\times 10^{13}$ G, at which the cyclotron energy equals the rest mass energy of an electron. A photon with energy $\epsilon$, moving at an angle $\theta_{\rm kB}$ relative to the local magnetic field B, can produce a pair when the condition $\epsilon^{2}\sin^{2}\theta_{\rm kB}/\bar{B}\gg 1$ is satisfied for pairs produced in high Landau states above threshold $2m_{e}c^{2}/\sin\theta_{\rm kB}$ (T. Erber, 1966; J. K. Daugherty & A. K. Harding, 1983) where $\bar{B}=B/B_{c}$.
In those cases the attenuation coefficient, ${\cal R}$, for pair production can be calculated using the expression (T. Erber, 1966; W.-Y. Tsai & T. Erber, 1974; A. K. Harding & D. Lai, 2006)
@@
where $l$ is the path length along the photon trajectory.
A photon with energy $\epsilon$ will escape the magnetosphere if its optical depth satisfies ${\cal T}(\epsilon,l)\ll 1$. A similar calculation as Eq. 27 has previously been considered for magnetospheric opacity of gamma-rays for isolated NSs; see e.g., A. K. Harding et al. (1997); M. G. Baring & A. K. Harding (2001); S. A. Story & M. G. Baring (2014); K. Hu et al. (2019, 2022) and references therein.
The photon escape energy can be proximity estimated by assuming that the mean path length is of the order of the lengthscale $c/\Omega$. Additionally, we consider average values for the magnetic field and the angle $\theta_{\rm kB}$ at the high-energy emission regions every time step. The attenuation coefficient was calculated from Eqs. (25)–(26) and then Eq. (27) calculated via integration for $l=c/\Omega$ for different $\epsilon$. The cutoff energy is considered the upper limit for $\epsilon$. Note that $\chi$, the attenuation coefficient $\cal{R}$, and consequently the optical depth and the escape energy vary over time as the orbital separation decreases due to the orbital inspiral. This evolution causes the outer magnetosphere to become increasingly opaque.
This behavior is evident in Figure 28, which shows the optical depth as a function of the photon energy for the case of $\alpha_{1}=0,\alpha_{2}=45,\lambda=1$ and $\Delta\Phi=0$, assuming stars with magnetic fields of $10^{12}$ G, left, and $10^{14}$ G, right. The line thickness represents the progression of time. The escape energy, defined as the energy where each line crosses the horizontal ${\cal T}=1$ threshold, decreases over time. For strong magnetic fields, the escaping energy decreases below the 1 MeV threshold (indicated by the vertical dashed line) required for pair production.
@@ 4 Discussion and Conclusions
The FF framework assumes ample plasma supply to screen accelerating electric field components while neglecting particle inertia, making it a well-suited approximation for studying large-scale magnetospheric EM interactions in the pre-merger phase. Our findings reveal strong Poynting flux emission anisotropies, which in turn induce forces on the two NSs, possibly affecting their orbital dynamics.
@@ 4.1 Global Magnetospheric Structure and Poynting Flux Evolution
Our results demonstrate that the interaction of the two stellar magnetospheres leads to a complex and dynamically evolving magnetic topology governed by the relative orientations and magnitudes of the magnetic moments.
The topology of the magnetosphere varies significantly with the alignment of the magnetic moments. In the aligned case ( $\uparrow$ $\uparrow$ ), field lines primarily bend toroidally, preventing direct connections between the two stars. In the anti-aligned case ( $\uparrow$ $\downarrow$ ), field lines link the two stars, forming flux tubes that evolve dynamically as the orbit tightens.
@@
The total Poynting flux exhibits strong anisotropies and varies significantly with the orbital phase. Compared to the aligned case, the anti-aligned configuration systematically produces higher flux, while cases with misaligned moments lead to periodic oscillations, with luminosity modulating at half the orbital period. These variations arise from the continuous restructuring of the magnetosphere and the evolution of polar caps along the inspiral. The luminosity reaches peak values of up to a few $10^{44}\,B_{12}^{2}\,r_{12\rm km}^{2}$ erg s-1. Notably, in the presence of a highly magnetized NS with $B\gtrsim 10^{15}$ G, Poynting luminosities can exceed $10^{50}$ erg s-1. Given the strong emission anisotropy, certain sight lines will experience isotropic-equivalent luminosities of $L_{\rm iso}\lesssim 10^{52}B_{15.5}^{2}/f_{\Omega,-1}$ for a beaming factor of $f_{\Omega,-1}\sim 0.1$ (e.g., see Figs. 7101215, and 16) and have been conjectured as the source of short GRBs (V. V. Usov, 1992). These results suggest that previous pessimistic evaluations regarding the detectability of pre-merger EM signals may need to be reconsidered.
@@
The degree of NS magnetization at merger is uncertain, as it depends on evolutionary factors such as field decay, accretion history, and the merger timescale. Accretion in recycled millisecond pulsars is known to suppress surface magnetic fields (e.g., G. S. Bisnovatyi-Kogan & B. V. Komberg, 1974; R. W. Romani, 1990; S. Konar & D. Bhattacharya, 1997; D. J. B. Payne & A. Melatos, 2004), and similar but weaker effects may occur in double NS systems, where limited accretion could reduce the field strength (T. M. Tauris et al., 2017; T. M. Tauris & E. P. J. van den Heuvel, 2023). However, recent studies suggest that some magnetars retain their strong magnetic fields for Myrs or longer (P. Beniamini et al., 2023), indicating that highly magnetized NSs may still be involved in BNS mergers. If such a magnetar participates in a merger, it could drive significant precursor EM radiation and introduce pronounced asymmetry in the magnetic moments of the two stars. The existence of “fast” but rare merger channels may also allow younger magnetars to merge before significant field decay occurs (P. Beniamini & T. Piran, 2024). While the overall prevalence of such systems remains uncertain (I. Mandel & F. S. Broekgaarden, 2022), multi-messenger observations may provide constraints on their frequency.
@@
Azimuthal phase offsets ($\Delta\Phi$) introduce time differences in flux evolution. As the orbital separation decreases, these offsets modify the final magnetic field geometry at the merger, potentially influencing post-merger magnetospheric activity and the resulting EM signals.
The scaling of EM luminosity with orbital frequency typically expressed as $\langle L\rangle\propto\Omega^{p}$, has been the subject of both theoretical and numerical studies, yielding a range of predictions that differ from $p=4$ for an isolated NS with a dipole field. Theoretical estimates for a magnetized star interacting with an unmagnetized companion suggest $p=13/3$ (D. Lai, 2012), assuming an energy dissipation model driven by flux tube reconnections. Models that account for shielding effects from a magnetized companion lead to a slightly shallower scaling of $p=10/3$ (C. Palenzuela et al., 2013a), while models considering the sum of the two luminosities to a scaling of $p=14/3$ (M. V. Medvedev & A. Loeb, 2013). Numerical simulations have reported a variety of power-law indices. Simulations tracking full inspirals found $p\approx 3/2$ (C. Palenzuela et al., 2013a) for equal dipole moments, deviating from their theoretically expected $p=10/3$. In highly asymmetric case ($\lambda=100$), the power law steepens to $p=14/3$ at earlier inspiral stages and reaches $p\approx 12$ just before the merger (C. Palenzuela et al., 2013a). Other simulations (E. R. Most & A. A. Philippov, 2020, 2022) using fixed orbital separations obtained $p\approx 7/3$ for peak luminosity scaling.
Our simulations reveal that the power-law index is not universal but depends sensitively on the magnetospheric configuration. Across our parameter survey, we find a wide range of exponents spanning $p\sim 1-6$ (Figure 17). However, regardless of inclination, we find a coarse scaling law of the average (over the different magnetic orientations) index $\overline{p}\sim(6\pm 0.2)-(3.6\pm 0.1)|\lambda|^{-0.21\pm 0.02}$, i.e. larger p is realized for larger $|\lambda|$ (Eq. 15).
@@
These results suggest that the pre-merger magnetospheric state plays a crucial role in shaping the overall evolution of EM luminosity. As discussed earlier, the variability in $p$ arises from the intrinsically complex nature of magnetospheric interactions, which themselves exhibit significant variation across the parameter space, leading to the continuous restructuring of open flux tubes throughout the inspiral. Our simulations confirm that these magnetospheric dynamics directly govern the evolution of open flux tubes, aligning with the observed luminosity variations.
4.2 Electromagnetically Induced Torques, Crustal Stresses, and Potential Gravitational Wave Imprints
@@
If magnetars are involved in such a BNS coalescence, this EM-induced deviation from a purely equatorial inspiral could imprint growing discrepancies in the gravitational waveform relative to standard unmagnetized NS merger models. Such systematic effects may be detectable with third-generation GW observatories, such as the Cosmic Explorer (B. P. Abbott et al., 2017b; M. Evans et al., 2021, 2023; A. Corsi et al., 2024) or the Einstein Telescope (M. Punturo et al., 2010; S. Hild et al., 2011; A. Abac et al., 2025). These instruments are expected to achieve exceptionally high signal-to-noise ratios, exceeding $10^{3}$ on a non-negligible subset of BNS mergers (S. Borhanian & B. S. Sathyaprakash, 2024), potentially revealing these previously unexplored magnetospheric effects on the GW signal.
However, while our post-processed analysis strongly suggests the presence of this accumulating angular momentum transfer, a fully self-consistent treatment is required to confirm its impact. Future simulations should incorporate the modified orbital dynamics caused by EM forces and evolve the system dynamically, allowing for feedback between the magnetosphere, the NS spins, and the binary’s motion and eccentricity. Such an approach will be crucial in assessing whether these effects leave indeed measurable imprints on GW signals (potentially hours before the merger) and determining their potential detectability.
Beyond influencing the binary’s orbital dynamics, the strong EM stresses induced by the evolving magnetosphere also act directly on the NS crusts, generating time-dependent, localized strain patterns. Our analysis reveals that these stresses vary over an orbital period, and for highly magnetized stars ($B\sim 10^{15}$ G), these stresses may become sufficiently strong to deform or even fracture the crust, potentially triggering magnetar-like bursting activity. If the crust fails, a transient X-ray burst similar to a magnetar giant flare could be produced, and the excitation of crustal normal modes could introduce additional observable signatures. These effects suggest that EM interactions in merging BNSs may not only modify the inspiral trajectory but could also lead to precursor high-energy transients detectable prior to coalescence.
@@ 4.3 High-Energy Signals
We also investigated the production of high-energy photons via curvature radiation from particles accelerated near the equatorial current sheets. We applied methodologies successfully used in pulsar studies, demonstrating how skymaps and light curves of high-energy emissions from BNSs can be constructed. These emission luminosities scale with the square root of the Poynting flux, following the heuristic dependence of $L_{\gamma}$ on the spin-down power at high $\dot{E}$ (D. A. Smith et al., 2023). Compared to the isolated pulsar case, we find a significantly more intricate emission structure, with light curves featuring extremely narrow peaks on the order of $<0.5$ ms. Detecting such rapid variability requires high temporal resolution and sensitivity, making Imaging Atmospheric Cherenkov Telescopes (IACTs) such as CTA strong candidates for observing these signals (Cherenkov Telescope Array Consortium et al., 2019; C. T. A. Observatory & C. T. A. Consortium, 2021). IACTs offer six to eight orders of magnitude better gamma-ray sensitivity (at higher energies) on short timescales compared to Fermi-LAT. However, advanced triplet-based photon analyses in Fermi-LAT might enable a statistically significant association (Fermi-LAT Collaboration et al., 2021; G. Principe et al., 2023) at lower photon energies. The periodic nature of these signals also allows for photon-based coherent search techniques (e.g., O. C. de Jager et al., 1989) hours or days before the merger, which can significantly enhance detectability. This requires precise pre-merger warnings and accurate localizations are available, capabilities that may only be feasible with third-generation GW networks or decihertz detectors.
@@
Our analysis indicates that B-$\gamma$ pair production is highly efficient in the strong magnetic fields present during the final inspiral stages. This leads to significant attenuation of high-energy photons, effectively limiting the maximum escaping energy to the MeV range in the last moments before the merger. As a result, while TeV to PeV emission is theoretically possible under radiation reaction limit assumptions, most escaping photons in the late inspiral phases are likely to be limited to $\lesssim 100$ MeV energies where competition between pair production and particle acceleration might approach a dynamic equilibrium. This suggests that the MeV band, where numerous instruments are currently selected or proposed (e.g., J. McEnery et al., 2019; R. Caputo et al., 2022; E. Orlando et al., 2022; J. Tomsick et al., 2024; T. Shutt et al., 2025), is the most promising range for detecting these emissions. Furthermore, as these pairs must be created in high Landau states, these pairs will subsequently radiate below 1 MeV, likely in the X-ray regime.
@@
Although magnetic pair production dominates attenuation, photon-photon ($\gamma\gamma$) pair production could also contribute, particularly in the earlier inspiral stages when the magnetosphere is more transparent. This process, owing to the threshold kinematics and cross-section, depends on the presence of a sufficient density of low-energy photons, which could arise from sources such as resonant crust shattering (D. Tsang et al., 2012; A. G. Suvorov et al., 2022), tidal heating or crust melt (A. J. Penner et al., 2012; P. Arras & N. N. Weinberg, 2019; Z. Pan et al., 2020; A. Passamonti et al., 2021), or pair synchrotron radiation. However, due to the strong suppression of high-energy photons by B-$\gamma$ attenuation, $\gamma\gamma$ pair production is expected to be secondary except possibly for $\sim 0.1-1$ GeV photons, which should interact efficiently with possibly extant local X-rays.
More critically, an open question remains regarding how much of the energy originally carried by high-energy photons is ultimately redistributed at lower frequencies by the pairs produced in both B-$\gamma$ and $\gamma\gamma$ interactions. Understanding the reprocessing efficiency of these pairs is crucial for assessing potential X-ray and optical emission, which may provide additional observational signatures beyond the MeV range.
@@
Unraveling the full complexity of these effects requires simulations that self-consistently incorporate the interplay between particle acceleration, pair production, and photon transport processes. The FF approximations used in this study provide valuable insights, but more detailed kinetic plasma simulations (e.g., particle-in-cell approaches) that incorporate detailed microphysics will be essential for fully resolving these open questions.
@@
A potential but uncertain source of precursor emission is coherent radio emission from these pair discharges, similar to polar cap radio generation in isolated pulsars (e.g., A. N. Timokhin, 2010; A. N. Timokhin & J. Arons, 2013; A. Philippov et al., 2020; J. Benáček et al., 2024). If such emission occurs, it could be detectable far beyond the GW horizon (e.g., A. J. Cooper et al., 2023). However, rate estimates suggest that BNS mergers contribute only a small fraction of apparently non-repeating FRBs (V. Ravi, 2019; B. Zhang, 2023; K. McGregor & D. R. Lorimer, 2024; B. Zhang, 2024), making this an unlikely but intriguing possibility.
4.4 Post-Merger Signals and Afterglow Echos

### v2

@@ 3.3.3 Pair Production and Gamma-ray Attenuation
This magnetic pair production is particularly significant in environments where the magnetic field strength approaches or exceeds the critical value, $B_{c}=4.414\times 10^{13}$ G, at which the cyclotron energy equals the rest mass energy of an electron. A photon with the energy $\epsilon$, moving at an angle $\theta_{\rm kB}$ relative to the local magnetic field B, can produce a pair when the condition $\epsilon^{2}\sin^{2}\theta_{\rm kB}/\bar{B}\gg 1$ is satisfied for pairs produced in high Landau states above the threshold $2m_{e}c^{2}/\sin\theta_{\rm kB}$ (T. Erber, 1966; J. K. Daugherty & A. K. Harding, 1983) where $\bar{B}=B/B_{c}$.
[…]
@@
${\cal T}(l)=\int_{0}^{l}{\cal R}\,ds,$ (28)
where $l$ is the path length along the photon trajectory.
A photon with the energy $\epsilon$ will escape the magnetosphere if its optical depth satisfies ${\cal T}(\epsilon,l)\ll 1$. A similar calculation as Eq. 28 has previously been considered for magnetospheric opacity of gamma-rays for isolated NSs; see, e.g., A. K. Harding et al. (1997),M. G. Baring & A. K. Harding (2001),S. A. Story & M. G. Baring (2014),K. Hu et al. (2019),K. Hu et al. (2022) and references therein.
@@
This behavior is evident in Figure 28, which shows the optical depth as a function of the photon energy for the case of $\alpha_{1}=0,\alpha_{2}=45,\lambda=1$, and $\Delta\Phi=0$, assuming stars with magnetic fields of $10^{12}$ G, left, and $10^{14}$ G, right. The line thickness represents the progression of time. The escape energy, defined as the energy where each line crosses the horizontal ${\cal T}=1$ threshold, decreases over time. For strong magnetic fields, the escaping energy decreases below the 1 MeV threshold (indicated by the vertical dashed line) required for pair production.
@@ 4 Discussion and Conclusions
The FF framework assumes ample plasma supply to screen accelerating electric field components while neglecting particle inertia, making it a well-suited approximation for studying large-scale magnetospheric EM interactions in the premerger phase. Our findings reveal strong Poynting flux emission anisotropies, which in turn induce forces on the two NSs, possibly affecting their orbital dynamics.
@@ 4.1 Global Magnetospheric Structure and Poynting Flux Evolution
Our results demonstrate that the interaction of the two stellar magnetospheres leads to a complex and dynamically evolving magnetic topology governed by the relative orientations and magnitudes of the magnetic moments. The overall structure of the magnetospheres and the distributions of the Poynting flux presented here exhibit similarities with other works presented in the literature (C. Palenzuela et al., 2013a; M. Ponce et al., 2014; E. R. Most & A. A. Philippov, 2022; J. F. Mahlmann & A. M. Beloborodov, 2025).
The topology of the magnetosphere varies significantly with the alignment of the magnetic moments. In the aligned case ( $\uparrow$ $\uparrow$ ), field lines primarily bend toroidally, preventing direct connections between the two stars. In the antialigned case ( $\uparrow$ $\downarrow$ ), field lines link the two stars, forming flux tubes that evolve dynamically as the orbit tightens. In both cases, the structure evolves with the time as expected from C. Palenzuela et al. (2013a) where the full Einstein-Maxwell-hydrodynamic equations are solved. Following the decrease of the separation between the two objects, the global structure shrinks as the merger time approaches.
@@
For the cases with $\lambda<0$ and $\alpha_{1}=0^{\circ}$, the magnetospheric structure exhibits flux tubes connecting the two stars, with topology determined by the orientation of their magnetic moments, as presented in the simulations of E. R. Most & A. A. Philippov (2022) where the stars are held at a constant separation. However, in Figures 1 and 3, presented in E. R. Most & A. A. Philippov (2022), the magnetic field lines become increasingly twisted, forming detached closed loops leading to a solar flare-like eruption. These features were postulated to potentially result in FRBs. In Figures 4, 8 and 13, for our cases whose configurations most closely match, we do not observe this behavior on the xzx-z plane (similar to C. Palenzuela et al., 2013a). However, we do see eruptive behavior off the xzx-z plane, as exhibited by time-dependent Poynting flux patterns in misaligned cases.
@@
The total Poynting flux exhibits strong anisotropies and varies significantly with the orbital phase. Compared to the aligned case, the antialigned configuration systematically produces higher flux, while cases with misaligned moments lead to periodic oscillations, with luminosity modulating at half the orbital period (M. Ponce et al., 2014; E. R. Most & A. A. Philippov, 2022; J. F. Mahlmann & A. M. Beloborodov, 2025). These variations arise from the continuous restructuring of the magnetosphere and the evolution of polar caps along the inspiral. The luminosity reaches peak values of up to a few $10^{44}\,B_{12}^{2}\,r_{12\rm km}^{2}$ erg s-1. Notably, in the presence of a highly magnetized NS with $B\gtrsim 10^{15}$ G, Poynting luminosities can exceed $10^{50}$ erg s-1. Given the strong emission anisotropy, certain sight lines will experience isotropic-equivalent luminosities of $L_{\rm iso}\lesssim 10^{52}B_{15.5}^{2}/f_{\Omega,-1}$ for a beaming factor of $f_{\Omega,-1}\sim 0.1$ (e.g., see Figs. 7101215, and 16) and have been conjectured as the source of short GRBs (V. V. Usov, 1992). These results suggest that previous pessimistic evaluations regarding the detectability of premerger EM signals may need to be reconsidered.
Although directly comparing the Poynting flux values presented here with those reported in other studies in the literature is difficult, as different NS parameters and assumptions were adopted, approximate scaling can be used to assess relative values. Furthermore, when comparing with simulations that assume constant separation, i.e., corresponding to earlier inspiral stages, one must account for the specific power-law dependence of the Poynting flux on orbital frequency (see subsection 3.1.5). Additionally, the simulations in other studies adopt different mass-to-radius ratios, leading to distinct orbital dynamics. As a result, discrepancies in the reported Poynting flux magnitude naturally arise. For example, when comparing the Poynting fluxes at the same time before merger, our results are higher by a factor of 2-3 for the antialigned case and 3.5-10 for the aligned case relative to the flux evolution reported in C. Palenzuela et al. (2013a), consistent with our higher initial orbital frequency. Conversely, our fluxes are lower by a factor 1.8-2.5 compared to the aligned cases in (J. F. Mahlmann & A. M. Beloborodov, 2025) as our simulations correspond to lower orbital frequency at their separation. Finally, relative to the antialigned cases presented in Figure 10 of E. R. Most & A. A. Philippov (2022), our fluxes are lower by roughly a factor of 10. We note, however, that the fluxes in that figure have been explicitly rescaled for visualization (by factors of 1/3 and 1/6), and it remains unclear whether further normalization or intercase scaling was applied, potentially contributing to the apparent discrepancy.
@@
The degree of NS magnetization at the merger is uncertain, as it depends on evolutionary factors such as field decay, accretion history, and the merger timescale. Accretion in recycled MSP is known to suppress surface magnetic fields (e.g., G. S. Bisnovatyi-Kogan & B. V. Komberg, 1974; R. W. Romani, 1990; S. Konar & D. Bhattacharya, 1997; D. J. B. Payne & A. Melatos, 2004), and similar but weaker effects may occur in double NS systems, where limited accretion could reduce the field strength (T. M. Tauris et al., 2017; T. M. Tauris & E. P. J. van den Heuvel, 2023). However, recent studies suggest that some magnetars retain their strong magnetic fields for Myr or longer (P. Beniamini et al., 2023), indicating that highly magnetized NSs may still be involved in BNS mergers. If such a magnetar participates in a merger, it could drive significant precursor EM radiation and introduce pronounced asymmetry in the magnetic moments of the two stars. The existence of “fast” but rare merger channels may also allow younger magnetars to merge before significant field decay occurs (P. Beniamini & T. Piran, 2024). While the overall prevalence of such systems remains uncertain (I. Mandel & F. S. Broekgaarden, 2022), multimessenger observations may provide constraints on their frequency.
@@
Azimuthal phase offsets ($\Delta\Phi$) introduce time differences in flux evolution. As the orbital separation decreases, these offsets modify the final magnetic field geometry at the merger, potentially influencing postmerger magnetospheric activity and the resulting EM signals.
The scaling of EM luminosity with orbital frequency typically expressed as $\langle L\rangle\propto\Omega^{p}$ has been the subject of both theoretical and numerical studies, yielding a range of predictions that differ from $p=4$ for an isolated NS with a dipole field. Theoretical estimates for a magnetized star interacting with an unmagnetized companion suggest $p=13/3$ (D. Lai, 2012), assuming an energy dissipation model driven by flux tube reconnections. Models that account for shielding effects from a magnetized companion lead to a slightly shallower scaling of $p=10/3$ (C. Palenzuela et al., 2013a), while models considering the sum of the two luminosities lead to a scaling of $p=14/3$ (M. V. Medvedev & A. Loeb, 2013). Numerical simulations have reported a variety of power-law indices. Simulations tracking full inspirals found $p\approx 3/2$ (C. Palenzuela et al., 2013a) for equal dipole moments, deviating from their theoretically expected $p=10/3$. In a highly asymmetric case ($\lambda=100$), the power law steepens to $p=14/3$ at earlier inspiral stages and reaches $p\approx 12$ just before the merger (C. Palenzuela et al., 2013a). Other simulations (E. R. Most & A. A. Philippov, 2020, 2022) using fixed orbital separations obtained $p\approx 7/3$ for peak luminosity scaling.
Our simulations reveal that the power-law index is not universal but depends sensitively on the magnetospheric configuration. Across our parameter survey, we find a wide range of exponents spanning $p\sim 1-6$ (Figure 17). However, regardless of inclination, we find a coarse scaling law of the average (over the different magnetic orientations) index $\overline{p}\sim(6\pm 0.2)-(3.6\pm 0.1)|\lambda|^{-0.21\pm 0.02}$, i.e., larger p is realized for larger $|\lambda|$ (Eq. 15).
@@
These results suggest that the premerger magnetospheric state plays a crucial role in shaping the overall evolution of EM luminosity. As discussed earlier, the variability in $p$ arises from the intrinsically complex nature of magnetospheric interactions, which themselves exhibit significant variation across the parameter space, leading to the continuous restructuring of open flux tubes throughout the inspiral. Our simulations confirm that these magnetospheric dynamics directly govern the evolution of open flux tubes, aligning with the observed luminosity variations.
4.2 Electromagnetically Induced Torques, Crustal Stresses, and Potential Gravitational-wave Imprints
@@
If magnetars are involved in such a BNS coalescence, this EM-induced deviation from a purely equatorial inspiral could imprint growing discrepancies in the gravitational waveform relative to standard unmagnetized NS merger models. Such systematic effects may be detectable with third-generation GW observatories, such as the Cosmic Explorer (B. P. Abbott et al., 2017b; M. Evans et al., 2021, 2023; A. Corsi et al., 2024) or the Einstein Telescope (M. Punturo et al., 2010; S. Hild et al., 2011; A. Abac et al., 2025). These instruments are expected to achieve exceptionally high signal-to-noise ratios, exceeding $10^{3}$ on a nonnegligible subset of BNS mergers (S. Borhanian & B. S. Sathyaprakash, 2024), potentially revealing these previously unexplored magnetospheric effects on the GW signal.
However, while our postprocessed analysis suggests the presence of this accumulating angular momentum transfer, a fully self-consistent treatment is required to confirm its impact. Future simulations should incorporate the modified orbital dynamics caused by EM forces and evolve the system dynamically, allowing for feedback between the magnetosphere, the NS spins, and the binary’s motion and eccentricity. Furthermore, the model of irrotational NSs inspiraling along quasi-circular orbits becomes increasingly approximate in the presence of strong magnetic interactions. As shown in K. Ioka & K. Taniguchi (2000), where the equations of motion are derived from the Euler-Lagrange equations, including the interaction term of static magnetic dipoles in a vacuum, circular orbit solutions may not exist when the magnetic moments are misaligned with the orbital axis. Although our simulations assume an FF magnetosphere rather than vacuum fields, these findings suggest that a more rigorous treatment of the orbital dynamics is necessary to accurately capture the backreaction effects of EM forces. Additionally, the approximation of spherical stars neglects magnetically induced deformations, i.e., ellipticities, which may be nonnegligible in highly magnetized systems (see S. Chandrasekhar, 1961; K. Ioka & K. Taniguchi, 2000, and references therein). Such an approach will be crucial in assessing whether these effects leave indeed measurable imprints on GW signals (potentially hours before the merger) and determining their potential detectability.
Beyond influencing the binary’s orbital dynamics, the strong EM stresses induced by the evolving magnetosphere also act directly on the NS crusts, generating time-dependent, localized strain patterns. Our analysis reveals that these stresses vary over an orbital period, and for highly magnetized stars ($B\sim 10^{15}$ G), these stresses may become sufficiently strong to deform or even fracture the crust, potentially triggering magnetar-like bursting activity. If the crust fails, a transient X-ray burst similar to a magnetar giant flare could be produced, and the excitation of crustal normal modes could introduce additional observable signatures. These effects suggest that EM interactions in merging BNSs not only may modify the inspiral trajectory but also could lead to precursor high-energy transients detectable prior to coalescence.
@@ 4.3 High-Energy Signals
We also investigated the production of high-energy photons via curvature radiation from particles accelerated near the equatorial current sheets. We applied methodologies successfully used in pulsar studies, demonstrating how skymaps and light curves of high-energy emissions from BNSs can be constructed. These emission luminosities scale with the square root of the Poynting flux, following the heuristic dependence of $L_{\gamma}$ on the spin-down power at high $\dot{E}$ (D. A. Smith et al., 2023). Compared to the isolated pulsar case, we find a significantly more intricate emission structure, with light curves featuring extremely narrow peaks on the order of <0.5<0.5 ms. Detecting such rapid variability requires high temporal resolution and sensitivity, making Imaging Atmospheric Cherenkov Telescopes (IACTs) such as Cherenkov Telescope Array (CTA) strong candidates for observing these signals (Cherenkov Telescope Array Consortium et al., 2019; C. T. A. Observatory & C. T. A. Consortium, 2021). IACTs offer 6 to 8 orders of magnitude better gamma-ray sensitivity (at higher energies) on short timescales compared to Fermi-LAT. However, advanced triplet-based photon analyses in Fermi-LAT might enable a statistically significant association (Fermi-LAT Collaboration et al., 2021; G. Principe et al., 2023) at lower photon energies. The periodic nature of these signals also allows for photon-based coherent search techniques (e.g., O. C. de Jager et al., 1989) hours or days before the merger, which can significantly enhance detectability. This requires precise premerger warnings and accurate localizations are available, capabilities that may only be feasible with third-generation GW networks or decihertz detectors.
@@
Our analysis indicates that B-$\gamma$ pair production is highly efficient in the strong magnetic fields present during the final inspiral stages. This leads to significant attenuation of high-energy photons, effectively limiting the maximum escaping energy to the MeV range in the last moments before the merger. As a result, while TeV to PeV emission is theoretically possible under radiation-reaction limit assumptions, most escaping photons in the late inspiral phases are likely to be limited to $\lesssim 100$ MeV energies where competition between pair production and particle acceleration might approach a dynamic equilibrium. This suggests that the MeV band, where numerous instruments are currently selected or proposed (e.g., J. McEnery et al., 2019; R. Caputo et al., 2022; E. Orlando et al., 2022; X. Pan et al., 2024; J. Tomsick et al., 2024; T. Shutt et al., 2025), is the most promising range for detecting these emissions. Furthermore, as these pairs must be created in high Landau states, these pairs will subsequently radiate below 1 MeV, likely in the X-ray regime.
@@
Although magnetic pair production dominates attenuation, photon-photon ($\gamma\gamma$) pair production could also contribute [^9] For instance, B. D. Metzger & C. Zivancev (2016) crudely estimated the compactness parameter and found it to be large., particularly in the earlier inspiral stages when the magnetosphere is more transparent. This process, owing to the threshold kinematics and cross section, however, depends on the presence of a sufficient density of low-energy photons, which could arise from sources such as resonant crust shattering (D. Tsang et al., 2012; A. G. Suvorov et al., 2022), tidal heating or crust melt (A. J. Penner et al., 2012; P. Arras & N. N. Weinberg, 2019; Z. Pan et al., 2020; A. Passamonti et al., 2021), or pair synchrotron radiation. However, due to the strong suppression of high-energy photons by B-$\gamma$ attenuation, $\gamma\gamma$ pair production is expected to be secondary except possibly for $\sim 0.1-1$ GeV photons, which should interact efficiently with possibly extant local X-rays. More critically, an open question remains regarding how much of the energy originally carried by high-energy photons is ultimately redistributed at lower frequencies by the pairs produced in both B-$\gamma$ and $\gamma\gamma$ interactions. Understanding the reprocessing efficiency of these pairs is crucial for assessing potential X-ray and optical emission, which may provide additional observational signatures beyond the MeV range.
@@
Unraveling the full complexity of these effects requires simulations that self-consistently incorporate the interplay between particle acceleration, pair production, and photon transport processes. The FF approximations used in this study provide valuable insights, but more detailed kinetic plasma simulations (e.g., PIC approaches) that incorporate detailed microphysics will be essential for fully resolving these open questions.
@@
A potential but uncertain source of precursor emission is coherent radio emission from these pair discharges, similar to polar cap radio generation in isolated pulsars (e.g., A. N. Timokhin, 2010; A. N. Timokhin & J. Arons, 2013; A. Philippov et al., 2020; J. Benáček et al., 2024). If such emission occurs, it could be detectable far beyond the GW horizon (e.g., A. J. Cooper et al., 2023). However, rate estimates suggest that BNS mergers contribute only a small fraction of apparently nonrepeating FRBs (V. Ravi, 2019; B. Zhang, 2023; K. McGregor & D. R. Lorimer, 2024; B. Zhang, 2024), making this an unlikely but intriguing possibility.
4.4 Postmerger Signals and Afterglow Echos